OpenStudy (anonymous):

HELP PLS THIS IS SERIOUS!!!!!!! D:

OpenStudy (jdoe0001):

3 ducks and 1 seagull with 5 snails

OpenStudy (anonymous):

can yall help meh find a website that simplifies an expression

OpenStudy (tylerd):

what kind of expression

OpenStudy (tylerd):

wolframalpha.com

OpenStudy (anonymous):

expression involving exponents

OpenStudy (tylerd):

it could be a log problem?

OpenStudy (jdoe0001):

hmmmm are you supposed to use a 'website that simplifies it for you" as opposed to do them youself?

OpenStudy (anonymous):

no its not ill give u an example: (6·7)^3

OpenStudy (tylerd):

if the seagul eats 1 duck how many snails are left?

OpenStudy (jdoe0001):

wait a sec..... seagulls do not eat ducks anyway =P

OpenStudy (anonymous):

its a joke lol

OpenStudy (jdoe0001):

\(\bf (6\cdot 7)^3\implies 6^3\cdot 7^3\)

OpenStudy (anonymous):

so can yall help meh find that website

OpenStudy (tylerd):

you can use a calculator for this

OpenStudy (jdoe0001):

hmmm what are you specifically after? are you supposed to use a website for that? you could just use a calculator I gather

OpenStudy (tylerd):

\[6^3 = 6\times6\times6\]

OpenStudy (anonymous):

let meh just show yall wut they want meh 2 do Simplify the expression if possible and write your answer as a power (using exponents). wen I put it in a calculator it comes up with numbers not exponents

OpenStudy (anonymous):

i need help on knowing how 2 change tha numbers 2 exponents

OpenStudy (tylerd):

ok evalute 6 times 7 first

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg <---- use that then, you should have covered that already anyhow

OpenStudy (anonymous):

i 4got how 2 evaluate :/ im sorry

OpenStudy (tylerd):

6 x 7 = 42 (42)^3

OpenStudy (tylerd):

its the same as 6^3 x 7^3

OpenStudy (anonymous):

oh ok thx Tyler :)